OpenStudy (anonymous):

difference quotient for the function f(x)= square root (7x+18)

OpenStudy (zzr0ck3r):

\[f(x) = \sqrt{7x+18}\\so\\\frac{\sqrt{7(x+h)+18}-\sqrt{7x+18}}{h}\]

OpenStudy (anonymous):

thank you! but how do i simplify that